OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (7, -3), (4, -8) i got: y=-8x-4

OpenStudy (anonymous):

hrmm. m=(-8--3)/(4-7)=-5/-3=5/3 y+3=(5/3)(x-7) =y=(5/3)x-(35/3)-3 =y=(5/3)x-(44/3) =(5/3)x-y=(44/3)

OpenStudy (anonymous):

didn't we do this one?

OpenStudy (anonymous):

im just confused on how to write the final answer

OpenStudy (anonymous):

m=(-3-(-8))/(7-4) m = 5/3

OpenStudy (anonymous):

m=(-3-(-8))/(7-4) m \[y-7=\frac{5}{3}(x-(-3)\] = 5/3

OpenStudy (anonymous):

write it like i wrote it in standard form

OpenStudy (anonymous):

\[y=\frac{5}{3}x+12\]

OpenStudy (anonymous):

now rewrite standard from

OpenStudy (anonymous):

none of your guys answers check out with those points

OpenStudy (anonymous):

\[-5x+3y=4\]

OpenStudy (anonymous):

=(5/3)x-y=(44/3)

OpenStudy (anonymous):

trust me

OpenStudy (anonymous):

I do this alot

OpenStudy (anonymous):

throw your equation into a calculator and throw in the x coordinated in table mode. it's wrong

OpenStudy (anonymous):

\[m=\frac{y _{2}-y _{1}}{x _{2}-x _{1}^{}}\]

OpenStudy (anonymous):

-3 is y2 -8 is y1

OpenStudy (anonymous):

repost the quetion and check some one answer